OpenStudy (anonymous):

what is the value of x if 4x=5y

OpenStudy (gorv):

x=5y/4

OpenStudy (anonymous):

no yar , option is 100, 110 120,150

OpenStudy (goformit100):

incomplete question.

OpenStudy (anonymous):

ur write, there is a figure in which a horizontal line and upon this horizontal line a inclined line, on both side of that inclined line written that x and y , now pls show method how will we solve,,,,,,,

OpenStudy (ajprincess):

It would be easy for us to help u if u attach the image:)

OpenStudy (anonymous):

ok i have attached athe image.

OpenStudy (anonymous):

well it need to add up to 180 no??

OpenStudy (anonymous):

bro ans will be 100 out of four question but how...

OpenStudy (ajprincess):

Assuming that they are angles x+y=180 4x=5y y=4x/5 x+4x/5=180 (5x+4x)/5=180 9x=180*5 9x=900 nw solve for x.:)

OpenStudy (anonymous):

really really thanks ...

OpenStudy (ajprincess):

u r welcome:)

OpenStudy (anonymous):

A horse is tied at the corner of a rectangular field, whose length is 20m and width is 16m with a rope whose length is 14 m. find the area which the horse can graze.

OpenStudy (ajprincess):

r there answer choices?

OpenStudy (anonymous):

154 sq. m. is ans but how?

OpenStudy (anonymous):

ok do.

OpenStudy (ajprincess):

since the horse is tied to a fixed point with the rope of length 14m the locus of it will be a circle with the radius 14m. the area it will cover within the field will be 1/4 of the area of the circle. area the horse grazes=pi*r^2/4 =(22/7)*(14*14)*(1/4)

OpenStudy (anonymous):

:- thanku again... ur too intelligent yar...

OpenStudy (ajprincess):

welcome and thnx for the compliment:)